# Rate-optimal and computationally efficient nonparametric estimation on the circle and the sphere

### v1

@@ 1 Introduction
Nonparametric Statistics introduces estimators of $f$, defined as functions of the form $\widehat{f}_{n}:\mathcal{M}^{n}\times\mathcal{M}\to\mathbb{R}$, where $\mathcal{M}$ denotes the set of all possible values that the random variable $X$ can take. Given the random sample $X_{1},\dots,X_{n}$, we can construct $\widehat{f}_{n}(x)=\widehat{f}_{n}(X_{1},\dots,X_{n};x)$, $x\in\mathcal{M}$, which is referred to as the density estimator of $f$.
@@
Historically, the earliest contributions are thanks to Rosenblatt [48], Parzen [42], and Breiman and Heller [4]. Foundational monographs include [50, 31], while the book [51] is now regarded as a standard modern reference. For a representative, but in no sense exhaustive, list of further developments in the field, see [1, 5, 6, 7, 11, 12, 17, 18, 19, 20, 21, 25, 26, 27, 28, 29, 30, 32, 33, 34, 37, 38, 39, 43, 44, 46, 47, 49].
In the present study we are interested in periodic random events and on random phenomena distributed on the Earth and Sky. The first can be modelled in terms of random variables distributed on the unit circle $\mathbb{S}^{1}=\{x\in\mathbb{R}^{2}:\|x\|=1\}$, while the second are expressed via random variables distributed on the unit sphere $\mathbb{S}^{2}=\{x\in\mathbb{R}^{3}:\|x\|=1\}$. Kernel density estimators (KDE) in that setting have a long history and applications in astrophysics; see for example [1], [2] and [22].
@@
The main challenge here is that the theoretically optimal estimators, established in the aforementioned works, are expressed in terms of infinite series. For example for the case of the sphere $\mathbb{S}^{d}$, a “smooth” density $f$, can be estimated by the KDE
@@
for some appropriate function $g:\mathbb{R}\to\mathbb{R}$, and some bandwidth $h>0$. Note that $P_{\ell}$ in (1.1) are the well-known Legendre polynomials; for more see Section 2.3. This automatically means that the KDE cannot be used by any software in its stand form. Therefore, when we are in front of a data-study, we need to address the next:
@@
The answer to the above question is affirmative. The first problem we address in this study is to establish that there exists an index $N$ such that the finite-order estimator, obtained by summing from $\ell=0$ to $\ell=N$, retains rate optimality. Note that this corresponds to a frequency cutoff, which is a natural practise in Approximation Theory. The corresponding results are provided in Section 3.
@@ 2.1 The underlying geometries
We proceed to the sphere $\mathbb{S}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:\|x\|=1\}$. We present the necessary notions released from [52, Section 6] and [16, Section 1]. the spherical coordinates are given by
@@
$\rho_{\mathbb{S}^{1}}(x,y)=\arccos(\langle x,y\rangle),\quad x,y\in\mathbb{S}^{2}.$
@@ 2.2 Legendre polynomials and spherical harmonics
Between the most valuable tools on the sphere, are the spherical harmonics and Legendre polynomials.
@@
Note further, that $P_{\ell}(1)=1$, for every $\ell\in\mathbb{N}_{0}$, $P_{0}\equiv 1$ and the set $\{P_{\ell}:\ell\in\mathbb{N}_{0}\}$ consist of an orthogonal basis for $L^{2}[-1,1]$, the space of square integrable functions on [1,1][-1,1].
@@
and are expanded for $\ell\leq m<0$ by the next conjugation;
@@ 2.3 Density estimation on the circle and sphere
Let $g:\mathbb{R}\to\mathbb{R}$ be a measurable function such that $g(0)=1$ with a “fast” decay (this is specified in Theorem 2.1). Such a function is referred as a symbol and will generate a KDE, as in the case of kernel functions on $\mathbb{R}^{d}$. Let $h>0$ a quantity that extends the notion of a bandwidth.
@@
$\widehat{f}_{n,h}(x)=\frac{1}{2\pi n}\sum_{j=1}^{n}\left(1+2\sum_{\ell=1}^{\infty}g(h\ell)\cos(\ell(\theta-\theta_{j}))\right),\quad x=x(\theta),\;\theta\in(-\pi,\pi],$ (2.8)
@@
$\widehat{f}_{n,h}(x)=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{\infty}\frac{2\ell+1}{4\pi}g\left(h\sqrt{(\ell(\ell+1))}\right)P_{\ell}(\langle x,X_{j}\rangle),\quad x\in\mathbb{S}^{2}.$ (2.9)
As always in Non-parametric statistics, we assume that the unknown density, lies on a regularity space. Let $s>0$. We say that a function $f$ belongs to the Hölder space of regularity order $s$, $\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$, when
@@
Observe that $L_{\mathbb{S}^{d}}$ are second order differential operators, which indicates their use for defining regularity spaces and the half in the exponent. For the fractional powers of the operators, we refer the reader to [45, p. 263]. For more details on regularity spaces on the sphere and more general spaces, we refer to [15, 24, 23, 35, 36]
@@
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g:\mathbb{R}\to\mathbb{R}$ be an even measurable function such that $g\in\mathcal{C}^{\tau}(\mathbb{R})$, for some $\tau>d+s$, $g(0)=1$, $g^{(\nu)}(0)=0$, for $1\leq\nu\leq\tau$, and for some $r>\tau+d$ and $C_{\tau}>0$,
$|g^{(\nu)}(\lambda)|\leq C_{\tau}(1+|\lambda|)^{-r},\quad\text{for every}\;\lambda\in\mathbb{R},\;\text{and}\;0\leq\nu\leq\tau.$
@@ 3 Rate-optimal and computationally efficient estimators
Whilst Theorem 2.1 shows that the KDEs in (2.8) and (2.9) are theoretically suitable estimators, the implementation of these KDEs comes with a computational challenge; each expression contains an infinite sum over $\ell$. To evaluate these estimators in practice, we must replace the infinite series by a finite-order version, keeping only terms up to some cutoff index $N$. Determining this index $N$ in such a way that the estimator remains rate-optimal motivates us to establish the next theorem, whose proof, as all proofs of our study, can be found in the Appendix:
@@
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g:\mathbb{R}\to\mathbb{R}$ be an even, measurable function such that $g\in\mathcal{C}^{\tau}(\mathbb{R})$, for some $\tau>d+s$, $g(0)=1$, $g^{(\nu)}(0)=0$, for $1\leq\nu\leq\tau$, and for some $r>\tau+d$ and $C_{\tau}>0$,
$|g^{(\nu)}(\lambda)|\leq C_{\tau}(1+|\lambda|)^{-r},\quad\text{for every}\;\lambda\in\mathbb{R},\;\text{and}\;0\leq\nu\leq\tau.$ (3.1)
We denote by
$N_{s}:=\left\lfloor\left(\dfrac{n^{(s+r)/(2s+d)}}{d\pi(r-d)}\right)^{1/(r-d)}\right\rfloor+1.$ (3.2)
@@
We close the present section by listing the next remarks.
@@
(a) Our finite-order KDES in (3.3) and (3.4) of Theorem 3.1, are computationally efficient, once we fix the appropriate symbol $g$. Let $d\in\{1,2\}$ and $s>0$ and set $r:=2d+\lceil s\rceil+1$. Then the symbol
@@
satisfies the assumptions of Theorem 3.1; see also [13, 14]. This means that our new optimal estimators are ready-to-go for any data analysis. For simulations and applications in the current study we refer the reader to the Sections 5 and 6.
(b) The finite estimators in Theorem 3.1, indeed integrate to the unit; $\int_{\mathbb{S}^{d}}\widehat{f}_{s}(x)d\sigma_{\mathbb{S}^{d}}(x)=1$, as one would expect. For the case of the circle $\mathbb{S}^{1}$, we can confirm this just by integrating (3.3) and using the properties of the trigonometric functions. For the case of the sphere $\mathbb{S}^{2}$, we just have to use the Funk–Hecke formula; [16, Theorem 1.2.9] to obtain for every $y\in\mathbb{S}^{2}$,
@@ 4 Probability estimation
A problem of fundamental interest is the estimation of the probability of an event occurring within a certain region of a space. Our kernel density estimators can be used for such an application, not only as plug-in estimators but also as a rate optimal estimation for such a probability on the circle or sphere. We introduce the following probability estimators for every measurable set $A\subset\mathbb{S}^{d}$, $d=1,2$ and we prove that these can indeed well-estimate the probability of any measurable subset $A$.
@@
Let $X$ a random variable distributed on $\mathbb{S}^{1}$, with density $f\in\mathcal{L}^{\infty}(\mathbb{S}^{1})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{1})$ for some $s>0$ and $X_{1},\dots,X_{n}$ iid samples of $X$. For any function $g:\mathbb{R}\to\mathbb{R}$ as in Theorem 3.1 and $h=n^{-1/(2s+1)}$. Then
@@
despite how simple form the domain $A$ has, and therefore the software suffers providing the final estimated probabilities.
@@
Note that for $u=1$, its value is equal to the standard Beta function. Furthermore, if $a,b\in\mathbb{N}$, then the integrand function in (4.4) is a polynomial, so the incomplete Beta function, its anti-derivative, is just a polynomial too. For the better presentation of the study we introduce the next kernels: For every $\ell\in\mathbb{N}$, $1\leq m\leq k\leq\ell$ and $\gamma_{1},\gamma_{2}\in[-1,1]$, we denote by
@@
Let $X$ a random variable distributed on $\mathbb{S}^{2}$, with density $f\in\mathcal{L}^{\infty}(\mathbb{S}^{2})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{2})$ for some $s>0$ and $X_{1},\dots,X_{n}$ iid samples of it. For any function $g:\mathbb{R}\to\mathbb{R}$ as in Theorem 3.1 and $h=n^{-1/(2s+2)}$. Then for every $\vartheta=(\vartheta_{1},\vartheta_{2}),\;\varphi=(\varphi_{1},\varphi_{2})$, with $0\leq\vartheta_{1}<\vartheta_{2}\leq\pi$ and $\pi\leq\varphi_{1}<\varphi_{2}\leq\pi$, we have that
@@ 5 Simulations
In this section, we reinforce with simulations the theoretical finds in Sections 3 and 4. We simulate from classic distributions on the sphere and circle and present our estimation of the densities and probabilities; that is the implementation of Theorem 3.1, Proposition 4.2 and Theorem 4.1. In particular, we first sample from the uniform distributions on $\mathbb{S}^{d}$ and then the well-known von Mises-Fisher distribution.
@@
For each simulated KDE, we estimated the probability over several regions using the methods established in Proposition 4.2 and Theorem 4.1 for $\mathbb{S}^{1}$ and $\mathbb{S}^{2}$, respectively. Precisely, the Tables 1-5 contain: these estimated probabilities compared to the observed frequency in the given region as well as the true (theoretical) probability, obtained by integrating the known PDF over the same region.
@@
$\text{MISE}(\widehat{f})=\mathbb{E}\left[\int_{\mathbb{S}^{d}}(\widehat{f}(x)-f(x))^{2}d\sigma_{\mathbb{S}^{d}}(x)\right],\quad d=1,2.$ (5.1)

### v2

# Rate-optimal and computationally efficient nonparametric density and probability estimation on the circle and the sphere

@@ 1 Introduction
Nonparametric Statistics introduces estimators of $f$, defined as functions of the form $\widehat{f}_{n}:\mathcal{M}^{n}\times\mathcal{M}\to\mathbb{R}$, where $\mathcal{M}$ denotes the set of all possible values that the random variable $X$ can take. Given the random sample $X_{1},\dots,X_{n}$, we can construct $\widehat{f}_{n}(x)=\widehat{f}_{n}(x;X_{1},\dots,X_{n})$, $x\in\mathcal{M}$, which is referred to as the density estimator of $f$.
@@
Historically, the earliest contributions are thanks to Rosenblatt [45], Parzen [40], and Breiman and Heller [4]. Foundational monographs include [27, 46], while the book [47] is now regarded as a standard modern reference. For a representative, but in no sense exhaustive, list of further developments in the field, see [5, 9, 10, 15, 16, 17, 18, 22, 23, 24, 25, 26, 28, 29, 30, 33, 34, 35, 41, 42, 44], and the references therein .
In the present study we are interested in periodic random events and in random phenomena distributed on the Earth and Sky. The first can be modelled in terms of random variables distributed on the unit circle $\mathbb{S}^{1}=\{x\in\mathbb{R}^{2}:\|x\|=1\}$, while the second are expressed via random variables distributed on the unit sphere $\mathbb{S}^{2}=\{x\in\mathbb{R}^{3}:\|x\|=1\}$. Kernel density estimators (KDE) in that setting have a long history and applications in astrophysics; see for example [1], [2] and [19].
@@
The main challenge here is that the theoretically optimal estimators, established in the aforementioned works, are expressed in terms of infinite series. For example for the case of the sphere $\mathbb{S}^{2}$, a “smooth” density $f$, can be estimated by the KDE
@@
for some appropriate function $g:\mathbb{R}\to\mathbb{R}$, and some bandwidth $h>0$. Note that $P_{\ell}$ in (1.1) are the well-known Legendre polynomials; for more see Section 2.3. This automatically means that the KDE cannot be used by any software in its stand form. Therefore, when we are in front of a data-study, we need to address the following:
@@
The answer to the above question is affirmative. The first problem we address in this study is to prove that there exists an index $N$ such that the finite-order estimator, obtained by summing from $\ell=0$ to $\ell=N$, retains rate optimality. Note that this corresponds to a frequency cutoff, which is a natural practise in Approximation Theory. The corresponding results are provided in Section 3.
@@ 2.1 The underlying geometries
We proceed to the sphere $\mathbb{S}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:\|x\|=1\}$. We present the necessary notions released from [48, Section 6] and [14, Section 1]. The spherical coordinates are given by
@@
$\rho_{\mathbb{S}^{2}}(x,y)=\arccos(\langle x,y\rangle),\quad x,y\in\mathbb{S}^{2}.$
@@ 2.2 Legendre polynomials and spherical harmonics
Among the most valuable tools on the sphere, are the spherical harmonics and Legendre polynomials.
@@
Note further, that $P_{\ell}(1)=1$, for every $\ell\in\mathbb{N}_{0}$, $P_{0}\equiv 1$ and the set $\{P_{\ell}:\ell\in\mathbb{N}_{0}\}$ consists of an orthogonal basis for $L^{2}[-1,1]$, the space of square integrable functions on [1,1][-1,1].
@@
and are expanded for $\ell\leq m<0$ by the following conjugation;
@@ 2.3 Density estimation on the circle and sphere
Let $g:\mathbb{R}\to\mathbb{R}$ be a measurable function such that $g(0)=1$ with a “fast” decay (this is specified in Theorem 2.1). Such a function is referred to as a symbol and will generate a KDE, as in the case of kernel functions on $\mathbb{R}^{d}$. Let $h>0$ be a quantity that extends the notion of a bandwidth.
@@
$\widehat{f}_{n,h}(x)=\frac{1}{2\pi n}\sum_{j=1}^{n}\Big(1+2\sum_{\ell=1}^{\infty}g(h\ell)\cos(\ell(\theta-\theta_{j}))\Big),\quad x=x(\theta),\;\theta\in(-\pi,\pi],$ (2.8)
@@
$\widehat{f}_{n,h}(x)=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{\infty}\frac{2\ell+1}{4\pi}g\Big(h\sqrt{(\ell(\ell+1))}\Big)P_{\ell}(\langle x,X_{j}\rangle),\quad x\in\mathbb{S}^{2}.$ (2.9)
As always in Nonparametric statistics, we assume that the unknown density, lies on a regularity space. Let $s>0$. We say that a function $f$ belongs to the Hölder space of regularity order $s$, $\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$, when
@@
Observe that $L_{\mathbb{S}^{d}}$ are second order differential operators, which indicates their use for defining regularity spaces and the half in the exponent. For the fractional powers of the operators, we refer the reader to [43, p. 263]. For more details on regularity spaces on the sphere and more general spaces, we refer to [13, 20, 21, 31, 32]
@@
We now fix some terminology that will be used throughout the paper. Let $s>0$. We say that a function $g:\mathbb{R}\to\mathbb{R}$ is an $s$-admissible symbol if it is even and measurable, belongs to $\mathcal{C}^{\tau}(\mathbb{R})$ for some $\tau>d+s$, satisfies $g(0)=1$ and $g^{(\nu)}(0)=0$ for all $1\leq\nu\leq\tau$, and if there exist constants $r>\tau+d$ and $C_{\tau}>0$ such that
$|g^{(\nu)}(\lambda)|\leq C_{\tau}(1+|\lambda|)^{-r},\qquad\text{for all }\lambda\in\mathbb{R},\;0\leq\nu\leq\tau.$ (2.11)
@@
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g$ and $s$-admissible symbol.
@@
The assumptions on the symbol are those required to generalize the usual notion of “kernels of order ss”, as given, for example, in [47, Definition 1.3], to general metric spaces. The details can be found in [12, Section 3].
@@ 3 Rate-optimal and computationally efficient estimators
Whilst Theorem 2.1 shows that the KDEs in (2.8) and (2.9) are theoretically suitable estimators, the implementation of these KDEs comes with a computational challenge; each expression contains an infinite sum over $\ell$. To evaluate these estimators in practise, we must replace the infinite series by a finite-order version, keeping only terms up to some cutoff index $N$. Determining this index $N$ in such a way that the estimator remains rate-optimal motivates us to establish the following theorem, whose proof, as all proofs of our study, can be found in the Appendix:
@@
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g$ and $s$-admissible symbol. We denote by
$N_{s}:=\Big\lfloor c_{rd}\,n^{\frac{s+r}{(2s+d)(r-d)}}\Big\rfloor+1,$ (3.1)
where $c_{rd}=(d\pi(r-d))^{-\frac{1}{r-d}}$.
@@
We close the present section by listing the following remarks.
@@
(a) Our finite-order KDES in (3.2) and (3.3) of Theorem 3.1, are computationally efficient, once we fix an appropriate $s$-admissible symbol $g$. Let $d\in\{1,2\}$ and $s>0$ and set $r:=2d+\lceil s\rceil+1$. Then the function
@@
is an $s$-admissible symbol [11, 12]. This means that our new optimal estimators are ready to be used in data analysis. This completes the line of work developed in [8, 11, 12], now complemented by the computational and applied aspects presented in the current study. For simulations and applications in the current study we refer the reader to Sections 5 and 6.
(b) The finite estimators in Theorem 3.1, indeed integrate to the unit; $\int_{\mathbb{S}^{d}}\widehat{f}_{s}(x)d\sigma_{\mathbb{S}^{d}}(x)=1$, as one would expect. For the case of the circle $\mathbb{S}^{1}$, we can confirm this just by integrating (3.2) and using the properties of the trigonometric functions. For the case of the sphere $\mathbb{S}^{2}$, we just have to use the Funk–Hecke formula; [14, Theorem 1.2.9] to obtain for every $y\in\mathbb{S}^{2}$,
@@
(d) For clarity, the stopping index $N_{s}$, defined in (3.1), depends on both the sample size $n$ and the regularity parameter $s$.
@@ 4 Probability estimation
A problem of fundamental interest is the estimation of the probability of an event occurring within a certain region of a space. Our kernel density estimators can be used for such an application, not only as plug-in estimators but also as a rate optimal estimation for such a probability on the circle or sphere. We introduce the following probability estimators for every measurable set $A\subset\mathbb{S}^{d}$, $d=1,2$ and we prove that these can indeed consistently estimate the probability of any measurable subset $A$.
@@
Let $X$ a random variable distributed on $\mathbb{S}^{1}$, with density $f\in\mathcal{L}^{\infty}(\mathbb{S}^{1})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{1})$ for some $s>0$ and $X_{1},\dots,X_{n}$ iid samples of $X$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g$ and $s$-admissible symbol. Then
@@
no matter how simple form the domain $A$ has, and therefore the software suffers providing the final estimated probabilities.
@@
Note that for $u=1$, its value is equal to the standard Beta function. Furthermore, if $a,b\in\mathbb{N}$, then the integrand function in (4.4) is a polynomial, so the incomplete Beta function, its anti-derivative, is just a polynomial too. For the better presentation of the study we introduce the following kernels: For every $\ell\in\mathbb{N}$, $1\leq m\leq k\leq\ell$ and $\gamma_{1},\gamma_{2}\in[-1,1]$, we denote by
@@
Let $X$ a random variable distributed on $\mathbb{S}^{2}$, with density $f\in\mathcal{L}^{\infty}(\mathbb{S}^{2})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{2})$ for some $s>0$ and $X_{1},\dots,X_{n}$ iid samples of it. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g$ and $s$-admissible symbol. Then for every $\vartheta=(\vartheta_{1},\vartheta_{2}),\;\varphi=(\varphi_{1},\varphi_{2})$, with $0\leq\vartheta_{1}<\vartheta_{2}\leq\pi$ and $\pi\leq\varphi_{1}<\varphi_{2}\leq\pi$, we have that
@@ 5 Simulations
In this section, we reinforce with simulations the theoretical finds in Sections 3 and 4. We simulate from classic distributions on the sphere and circle and present our estimation of the densities and probabilities; that is the implementation of Theorem 3.1, Proposition 4.2 and Theorem 4.1. In particular, we first sample from the uniform distributions on $\mathbb{S}^{d}$ and then the very well-known von Mises-Fisher distribution.
@@
For each simulated KDE, we estimated the probability over several regions using the methods established in Proposition 4.2 and Theorem 4.1 for $\mathbb{S}^{1}$ and $\mathbb{S}^{2}$, respectively. Precisely, the Tables 1-6 contain: these estimated probabilities compared to the observed frequency in the given region as well as the true (theoretical) probability, obtained by integrating the known PDF over the same region.
@@
$\text{MISE}(\widehat{f})=\mathbb{E}_{f}\Big[\int_{\mathbb{S}^{d}}(\widehat{f}(x)-f(x))^{2}d\sigma_{\mathbb{S}^{d}}(x)\Big],\quad d=1,2.$ (5.1)